Consider a computer system in which a physical page number is 20 bits, a virtual page number is 50 bits,

and a virtual address is 64 bits.

What is the maximum amount of physical memory, in GiB, that this system could have?
Answer #1

virtual page number bits = virtual address - offset

50 = 64 - offset

Therefore offset = 64 - 50 = 14

physical page number bits = physical address - offset

Therefore physical address = physical page number bits + offset

= 20 + 14 = 34

Hence physical address has 34 bits.

From this maximum amount of physical memory = 17.17 GB(2^32 bytes)